# Weakly-Supervised Conditional Embedding for Referred Visual Search

### v2

@@ Abstract
This paper introduces a new challenge for image similarity search in the context of fashion, addressing the inherent ambiguity in this domain stemming from complex images. We present Referred Visual Search (RVS), a task allowing users to define more precisely the desired similarity, following recent interest in the industry. We release a new large public dataset, LAION-RVS-Fashion, consisting of 272k fashion products with 842k images extracted from LAION, designed explicitly for this task. However, unlike traditional visual search methods in the industry, we demonstrate that superior performance can be achieved by bypassing explicit object detection and adopting weakly-supervised conditional contrastive learning on image tuples. Our method is lightweight and demonstrates robustness, reaching Recall at one superior to strong detection-based baselines against 2M distractors. [^1] Code, data and models are available at https://github.com/Simon-Lepage/CondViT-LRVSF
Figure 1: Overview of the Referred Visual Search task. Given a query image and conditioning information, the goal is to retrieve a target instance from a large scale collection. Note that a query is made of one image plus an additional text or category, precising what aspect of the image is relevant for the query.
@@ 1 Introduction
Image embeddings generated by deep neural networks play a crucial role in a wide range of computer vision tasks. Image retrieval, has gained substantial prominence, leading to the development of dedicated vector database systems (Johnson et al., 2019). These systems facilitate efficient retrieval by comparing embedding values and identifying the most similar images within the database.
@@
An intuitive approach is to request users furnish supplementary information delineating their interests, such as providing an image of an individual and denoting interest in the hat (see Fig. 1). Numerous industry leaders including Google, Amazon, and Pinterest have adopted this tactic, however academic discourse on potential alternative methodologies for this task remains scarce. For convenience, we propose terming this task Referred Visual Search (RVS), as it is likely to garner attention from the computer vision community due to the utility for product search in extensive catalogs.
In practice, object selection in complex scenes is classically tackled using object detection and crops (Jing et al., 2015; Hu et al., 2018; Ge et al., 2019; Shiau et al., 2020). Some recent approaches use categorical attributes (Dong et al., 2021) or text instead (Das et al., 2022), and automatically crop the image based on learned attention to input attributes. Indeed it is also possible to ask the user to perform the crop himself, yet in all the situations the performance of the retrieval will be sensitive to this extraction step making it costly to build a generic retrieval tool. Recently, Jiao et al. (2023) went a step further, incorporating prior knowledge about the taxonomy of fashion attributes and classes without using crops. They use a multi-granularity loss and two sub-networks to learn attribute and class-specific representations, resulting in improved robustness for fashion retrieval, yet without providing any code. In this work, we seek to advance in this direction and totally eliminate the need for explicit detection or segmentation while still producing similarities in the embedding space specific to the conditioning. Indeed, all localization-dependent approaches hinge on multi-stage processing heuristics particular to the dataset, whereas an end-to-end approach has the potential to be more generalizable and robust. To our knowledge, no public dataset is available for this task. Therefore, to demonstrate our approach’s soundness, we extracted a subset of LAION 5B focused on pairs of images sharing a labeled similarity in the domain of Fashion.
@@
  • The introduction of a new dataset, referred to as LAION-RVS-Fashion, which is derived from the LAION-5B dataset and comprises 272k fashion products with nearly 842k images. This dataset features a test set with an addition of more than 2M distractors, enabling the evaluation of method robustness in relation to gallery size. The dataset’s pairs and additional metadata are designed to necessitate the extraction of particular features from complex images.
@@ 2 Related Work
Multi-modal Models.
Deep learning has made significant progress in both vision and language domains, leading to the emergence of new vision-language methods. This new field notably developed Vision-Language Pre-Training (Du et al., 2022b), leveraging many pretext tasks to create models that can be finetuned for downstream multi-modal applications. Specific models have been trained on fashion datasets to extract more relevant features (Zhuge et al., 2021; Mirchandani et al., 2022; Goenka et al., 2022; Ji et al., 2023), and applied to multi-modal product retrieval (Zhan et al., 2021; Yu et al., 2022; Zheng et al., 2023). In our work, we use CLIP (Radford et al., 2021) as a general feature extractor for our visual encoder.
Vision-Language processing also brought new challenges, in particular Referring Expression Comprehension and Segmentation where a sentence designates an object in a scene, that the network has to localize. For the comprehension task (similar to open-vocabulary object detection) the goal is to output a bounding box (Luo et al., 2020; Zeng et al., 2022a; b; Liu et al., 2023). The segmentation task aims at producing an instance mask for images (Zhang et al., 2017; Luo et al., 2020; Huang et al., 2020; Ding et al., 2021; Kirillov et al., 2023) and recently videos (Wu et al., 2022; Botach et al., 2022). In this paper, we propose a referring expression task, where the goal is to embed the designated object of interest into a representation that can be used for retrieval. We explore the use of Grounding DINO (Liu et al., 2023) and Segment Anything (Kirillov et al., 2023) to create a strong baseline on our task.
@@
In the last decade, content-based image retrieval has changed because of the arrival of deep learning, which replaced many handcrafted heuristics (keypoint extraction, descriptors, geometric matching, re-ranking…) (Dubey, 2022). In the industry this technology has been of interest to retail companies and search engines to develop visual search solutions, with new challenges stemming from the large scale of such databases. Initially using generic pretrained backbones to extract embeddings with minimal retraining (Yang et al., 2017), methods have evolved toward domain-specific embeddings supervised by semantic labels, and then multi-task domain-specific embeddings, leveraging additional product informations (Zhai et al., 2019; Bell et al., 2020; Tran et al., 2019). The latest developments in the field incorporate multi-modal features for text-image matching (Zhan et al., 2021; Yu et al., 2022; Zheng et al., 2023), with specific vision-language pretext tasks.
However, these methods often consider that the query image is unambiguous, and often rely on a region proposal system to crop the initial image (Jing et al., 2015; Zhang et al., 2018; Hu et al., 2018; Shiau et al., 2020; Bell et al., 2020; Du et al., 2022a). In our work, we bypass this step and propose an end-to-end framework, leveraging the Transformer architecture to implicitly perform this detection step conditionally to the referring information.
@@
Conditional similarity search has been attempted in two ways: by modifying the retrieval process or the embedding process. On one hand, for the retrieval process, Hamilton et al. (2021) propose to use a dynamically pruned random projection tree. On the other hand, for the embeddings, previous work in conditional visual similarity learning has been oriented toward attribute-specific retrieval, considering that different similarity spaces should be defined depending on chosen discriminative attributes (Veit et al., 2017; Mu & Guttag, 2022). These approaches use either a mask applied on the features (Veit et al., 2017), or different projection heads (Mu & Guttag, 2022), and require extensive data labeling.
@@
In Fashion, ASEN (Ma et al., 2020) uses spatial and channel attention to an attribute embedding to extract specific features in a global branch. Dong et al. (2021) and Das et al. (2022) build upon this model and add a local branch working on an attention-based crop. Recently, Jiao et al. (2023) incorporated prior knowledge about fashion taxonomy in this process to create class-conditional embeddings based on known fine-grained attributes, using multiple attribute-conditional attention modules. In a different domain, Asai et al. (2022) tackle a conditional document retrieval task, where the user intent is made explicit by concatenating instructions to the query documents. In our work, we use Vision Transformers (Dosovitskiy et al., 2021) to implicitly pool features depending on the conditioning information, without relying on explicit ROI cropping or labeled fine-grained attributes.
Composed Image Retrieval.
@@
Composed Image Retrieval (CIR) (Vo et al., 2019) is another retrieval task where the embedding of an image must be modified following a given instruction. Recent methods use a composer network after embedding the image and the modifying text (Lee et al., 2021; Chen et al., 2022; Baldrati et al., 2023). While CIR shares similarities with RVS in terms of inputs and outputs, it differs conceptually. Our task focuses on retrieving items based on depicted attributes and specifying a similarity computation method, rather than modifying the image.
In Fashion, CIR has been extended to dialog-based interactive retrieval, where an image query is iteratively refined following user instructions (Guo et al., 2018; Wu et al., 2019; Yuan & Lam, 2021; Han et al., 2022).
Retrieval Datasets.
Standard datasets in metric learning literature consider that the images are object-centric, and focus on single salient objects (Wah et al., 2011; Krause et al., 2013; Song et al., 2016). In the fashion domain there exist multiple datasets dedicated to product retrieval, with paired images depicting the same product and additional labeled attributes. A recurrent focus of such datasets is cross-domain retrieval, where the goal is to retrieve images of a given product taken in different situations, for exemple consumer-to-shop (Liu et al., 2012; Wang et al., 2016; Liu et al., 2016; Ge et al., 2019), or studio-to-shop (Liu et al., 2016; Lasserre et al., 2018). The domain gap is in itself a challenge, with issues stemming from irregular lighting, occlusions, viewpoints, or distracting backgrounds. However, the query domain (consumer images for exemple) often contains scenes with multiple objects, making queries ambiguous. This issue has been circumvented with the use of object detectors and landmarks detectors (Kiapour et al., 2015; Huang et al., 2015; Liu et al., 2016; Ge et al., 2019). (Kiapour et al., 2015; Liu et al., 2016; Wang et al., 2016) are not accessible anymore.
With more than 272k distinct training product identities captured in multi-instance scenes, our new dataset proposes an exact matching task similar to the private Zalando dataset (Lasserre et al., 2018), while being larger than existing fashion retrieval datasets and publicly available. We also create an opportunity for new multi-modal approaches, with captions referring to the product of interest in each complex image, and for robustness to gallery size with 2M added distractors at test time.
@@ 3 Conditional Embedding
Figure 2: Overview of our method on LRVS-F. For each element in a batch, we embed the scene conditionally and the isolated item unconditionally. We optimize a Normalized Temperature-scaled Cross Entropy loss over the cosine similarity matrix. $\oplus$ denotes concatenation to the patch sequence.
Task Formulation:
@@
We implement $\phi$ by modifying the Vision Transformer (ViT) architecture (Dosovitskiy et al., 2021). The conditioning is an additional input token with an associated learnable positional encoding, concatenated to the sequence of image patches. The content of this token can either be learned directly (e.g. for discrete categorical conditioning), or be generated by another network (e.g. for textual conditioning). At the end of the network, we linearly project the [CLS] token to map the features to a metric space. We experimented with concatenating at different layers in the transformer, and found that concatenating before the first layer is the most sensible choice (see Appendix B.3).
We train the network with the InfoNCE loss (Sohn, 2016; Oord et al., 2018), following CLIP (Radford et al., 2021), which is detailed in the next paragraph. However, we hypothesize that even though our method relies on a contrastive loss, it does not explicitly require a specific formulation of it. We choose the InfoNCE loss because of its popularity and scalability.
During training, given a batch of $N$ pairs of images and conditioning $((x_{i}^{A},c_{i}^{A});(x_{i}^{B},c_{i}^{B}))_{i=1..N}$, we compute their conditional embeddings $(z_{i}^{A},z_{i}^{B})_{i=1..N}$ with $z=\phi(x,c)\in\mathbb{R}^{d}$. We compute a similarity matrix $S$ where $S_{ij}=s(z_{i}^{A},z_{j}^{B})$, with $s$ the cosine similarity. We then optimize the similarity of the correct pair with a cross-entropy loss, effectively considering the $N-1$ other products in the batch as negatives:
@@
with $\tau$ a learned temperature parameter, and the final loss is $\mathcal{L}=l(S)/2+l(S^{\top})/2$. Please refer to Fig. 2 for an overview of the method. The $\tau$ parameter is used to follow the initial formulation of CLIP (Radford et al., 2021) and is optimized by gradient during the training.
At test time, we use FAISS (Johnson et al., 2019) to create a unique index for the entire gallery and perform fast similarity search on GPUs.
4 Dataset
Metric learning methods work by extracting features that pull together images labeled as similar (Dubey, 2022). In our case, we wanted to create a dataset where this embedding has to focus on a specific object in a scene to succeed. We found such images in fashion, thanks to a standard practice in this field consisting in taking pictures of the products alone on neutral backgrounds, and worn by models in scenes involving other clothing items (see Fig. 3).
We created LAION-RVS-Fashion (abbreviated LRVS-F) from LAION-5B by collecting images of products isolated and in context, which we respectively call simple and complex. We grouped them using extracted product identifiers. We also gathered and created a set of metadata to be used as referring information, namely LAION captions, generated captions, and generated item categories. Please refer to Appendix A.2, A.3 and C for metadata details and a datasheet (Gebru et al., 2021).
In total, we extracted 272,451 products for training, represented in 841,718 images. This represents 581,526 potential simple/complex positive pairs. We additionally extracted 400 products (800 images) to create a validation set, and 2,000 products (4,000 images) for a test set. We added 99,541 simple images in the validation gallery as distractors, and 2,000,014 in the test gallery. Details about the benchmark will be given in section 4.2.
Product 1 Product 2
Refer to captionRefer to captionRefer to caption Refer to captionRefer to caption
Category Lower Body Bags
LAION Text Michael Kors ’Samantha’ skinny trousers BARK - striped tote 7
BLIP2 Caption a women’s beige trousers a handbag with navy and white stripes
4.1 Construction
Image Collection.
Annotations.
We generated synthetic labels for the image complexity, the category of the product, and added new captions to replace the noisy LAION alt-texts. For the complexity labels, we employed active learning to incrementally train a classifier to discern between isolated objects on neutral backdrops and photoshoot scenes. The product categories were formed by aggregating various fine-grained apparel items into 10 coarse groupings. This categorization followed the same active learning protocol. Furthermore, the original LAION captions exhibited excessive noise, including partial translations or raw product identifiers. Therefore, we utilized BLIP-2 (Li et al., 2023) to generate new, more descriptive captions.
Dataset Cleaning.
In order to mitigate false negatives in our results, we utilized Locally Sensitive Hashing and OpenCLIP ViT-B/16 embeddings to eliminate duplicates. Specifically, we removed duplicates between the test targets and test distractors, as well as between the validation targets and validation distractors. Throughout our experiments, we did not observe any false negatives in the results. However, there remains a small quantity of near-duplicates among the distractor images.
4.2 Benchmark
For this dataset, we propose to frame the benchmark as an asymmetric task : the representation of simple images (the gallery) should not be computed conditionally. This choice is motivated by three reasons. First, when using precise free-form conditioning (such as LAION texts, which contain hashed product identifiers and product names) a symmetric encoding would enable a retrieval based solely on this information, completely disregarding the image query. Second, for discrete (categorical) conditioning it allows the presence of items of unknown category in the gallery, which is a situation that may occur in distractors. Third, these images only depict a single object, thus making referring information unnecessary. A similar setting is used by Asai et al. (2022).

### v3

# LRVS-Fashion: Extending Visual Search with Referring Instructions

@@ Abstract
This paper introduces a new challenge for image similarity search in the context of fashion, addressing the inherent ambiguity in this domain stemming from complex images. We present Referred Visual Search (RVS), a task allowing users to define more precisely the desired similarity, following recent interest in the industry. We release a new large public dataset, LRVS-Fashion, consisting of 272k fashion products with 842k images extracted from fashion catalogs, designed explicitly for this task. However, unlike traditional visual search methods in the industry, we demonstrate that superior performance can be achieved by bypassing explicit object detection and adopting weakly-supervised conditional contrastive learning on image tuples. Our method is lightweight and demonstrates robustness, reaching Recall at one superior to strong detection-based baselines against 2M distractors. [^1] The dataset is available at https://huggingface.co/datasets/Slep/LAION-RVS-Fashion
Figure 1: Overview of the Referred Visual Search task. Given a query image and conditioning information, the goal is to retrieve a target instance from a large gallery. Note that a query is made of an image and an additional text or category, precising what aspect of the image is relevant.
@@ 1 Introduction
Image embeddings generated by deep neural networks play a crucial role in a wide range of computer vision tasks. Image retrieval has gained substantial prominence, leading to the development of dedicated vector database systems [22]. These systems facilitate efficient retrieval by comparing embedding values and identifying the most similar images within the database.
@@
An intuitive approach is to request users furnish supplementary information delineating their interests, such as providing an image of an individual and denoting interest in the hat (see Fig. 1). Numerous industry leaders including Google, Amazon, and Pinterest have adopted this tactic, however academic discourse on potential alternative methodologies for this task remains scarce as the domain lacks dedicated datasets. For convenience, we propose terming this task Referred Visual Search (RVS), as it is likely to garner attention from the computer vision community due to the utility for product search in extensive catalogs.
In practice, object selection in complex scenes is classically tackled using object detection and crops [21, 17, 12, 42]. Some recent approaches use categorical attributes [8] or text instead [6], and automatically crop the image based on learned attention to input attributes. It is also possible to ask the user to perform the crop himself, yet in all the situations the performance of the retrieval will be sensitive to this extraction step making it costly to build a generic retrieval tool. Recently, Jiao et al. [20] went a step further, incorporating prior knowledge about the taxonomy of fashion attributes and classes without using crops. They use a multi-granularity loss and two sub-networks to learn attribute and class-specific representations, resulting in improved robustness for fashion retrieval, yet without providing any code.
In this work, we seek to support these efforts by providing a dataset dedicated to RVS. We extracted a subset of LAION 5B [41] focused on pairs of images sharing a labeled similarity in the domain of fashion, and propose a method to eliminate the need for explicit detection or segmentation, while still producing similarities in the embedding space specific to the conditioning. We think that such end-to-end approach has the potential to be more generalizable and robust, whereas localization-dependent approaches hinge on multi-stage processing heuristics specific to the dataset.
@@
  • The introduction of a new dataset, referred to as LRVS-Fashion, which is derived from the LAION-5B dataset and comprises 272k fashion products with nearly 842k images. This dataset features a test set with an addition of more than 2M distractors, enabling the evaluation of method robustness in relation to gallery size. The dataset’s pairs and additional metadata are designed to necessitate the extraction of particular features from complex images.
@@ 2 Related Work
Retrieval Datasets.
Standard datasets in metric learning literature consider that the images are object-centric, and focus on single salient objects [49, 25, 45]. In the fashion domain there exist multiple datasets dedicated to product retrieval, with paired images depicting the same product and additional labeled attributes. A recurrent focus of such datasets is cross-domain retrieval, where the goal is to retrieve images of a given product taken in different situations, for exemple consumer-to-shop [31, 50, 32, 12], or studio-to-shop [32, 27]. The domain gap is in itself a challenge, with issues stemming from irregular lighting, occlusions, viewpoints, or distracting backgrounds. However, the query domain (consumer images for exemple) often contains scenes with multiple objects, making queries ambiguous. This issue has been circumvented with the use of object detectors and landmarks detectors [23, 18, 32, 12]. Some are not accessible anymore [23, 32, 50].
With more than 272k distinct training product identities captured in multi-instance scenes, our new dataset proposes an exact matching task similar to the private Zalando dataset [27], while being larger than existing fashion retrieval datasets and publicly available. We also create an opportunity for new multi-modal approaches, with captions referring to the product of interest in each complex image, and for robustness to gallery size with 2M added distractors at test time.
@@
In the last decade, content-based image retrieval has changed because of the arrival of deep learning, which replaced many handcrafted heuristics (keypoint extraction, descriptors, geometric matching, re-ranking…) [11]. In the industry this technology has been of interest to retail companies and search engines to develop visual search solutions, with new challenges stemming from the large scale of such databases. Initially using generic pretrained backbones to extract embeddings with minimal retraining [53], methods have evolved toward domain-specific embeddings supervised by semantic labels, and then multi-task domain-specific embeddings, leveraging additional product informations [58, 3, 46]. The latest developments in the field incorporate multi-modal features for text-image matching [59, 54, 62], with specific vision-language pretext tasks.
However, these methods often consider that the query image is unambiguous, and often rely on a region proposal system to crop the initial image [21, 60, 17, 42, 3, 10]. In our work, we bypass this step and propose an end-to-end framework, leveraging the Transformer architecture to implicitly perform this detection step conditionally to the referring information.
Referring Tasks.
Referring tasks are popular in vision-language processing, in particular Referring Expression Comprehension and Segmentation where a sentence designates an object in a scene, that the network has to localize. For the comprehension task (similar to open-vocabulary object detection) the goal is to output a bounding box [34, 56, 57, 30]. The segmentation task aims at producing an instance mask for images [61, 34, 19, 7, 24] and recently videos [52, 4]. In this paper, we propose a referring expression task, where the goal is to embed the designated object of interest into a representation that can be used for retrieval. We explore the use of Grounding DINO [30] and Segment Anything [24] to create a strong baseline on our task.
@@
Conditional similarity search has been studied through the retrieval process and the embedding process. On one hand, for the retrieval process, Hamilton et al. [15] propose to use a dynamically pruned random projection tree. On the other hand, previous work in conditional visual similarity learning focused on attribute-specific retrieval, defining different similarity spaces depending on chosen discriminative attributes [47, 36]. They use either a mask applied on the features [47], or different projection heads [36], and require extensive data labeling.
@@
In Fashion, ASEN [35] uses spatial and channel attention to an attribute embedding to extract specific features in a global branch. Dong et al. [8] and Das et al. [6] build upon this model and add a local branch working on an attention-based crop. Recently, Jiao et al. [20] incorporated prior knowledge about fashion taxonomy in this process to create class-conditional embeddings based on known fine-grained attributes, using multiple attribute-conditional attention modules. In a different domain, Asai et al. [1] tackle a conditional document retrieval task, where the user intent is made explicit by concatenating instructions to the query documents. In our work, we use Vision Transformers [9] to implicitly pool features depending on the conditioning information, without relying on explicit ROI cropping or labeled fine-grained attributes.
@@
Composed Image Retrieval (CIR) [48] is another retrieval task where the embedding of an image must be modified following a given instruction. Recent methods use a composer network after embedding the image and the modifying text [28, 5, 2]. While CIR shares similarities with RVS in terms of inputs and outputs, it differs conceptually. Our task focuses on retrieving items based on depicted attributes and specifying a similarity computation method, rather than modifying the image. In Fashion, CIR has been extended to dialog-based interactive retrieval, where an image query is iteratively refined following user instructions [14, 51, 55, 16].
Figure 2: Overview of the data collection. a) Selection of a subset of domains belonging to known fashion retailers. b) Extraction of product identifiers in the URLs using domain-specific regular expressions. c) Generation of synthetic metadata for the products (categories, captions, …) using both pretrained and finetuned models. d) Deduplication of the images, and assignment to subsets.
3 Dataset
Metric learning methods work by extracting features that pull together images labeled as similar [11]. In our case, we wanted to create a dataset where this embedding has to focus on a specific object in a scene to succeed. We found such images in fashion, thanks to a standard practice in this field consisting in taking pictures of the products alone on neutral backgrounds, and worn by models in scenes involving other clothing items (see Fig. 3).
We created LAION-RVS-Fashion (abbreviated LRVS-F) from LAION-5B by collecting images of products isolated and in context, which we respectively call simple and complex. We grouped them using extracted product identifiers. We also gathered and created a set of metadata to be used as referring information, namely LAION captions, generated captions, and generated item categories. The process is depicted Fig. 2, presented in Section 3.1 with additional details in Appendix A.3.
3.1 Construction
Image Collection.
Annotations.
We generated synthetic labels for the image complexity, the category of the product, and added new captions to replace the noisy LAION alt-texts. For the complexity labels, we employed active learning to incrementally train a classifier to discern between isolated objects on neutral backdrops and photoshoot scenes. The product categories were formed by aggregating various fine-grained apparel items into 10 coarse groupings. This categorization followed the same active learning protocol. Furthermore, the original LAION captions exhibited excessive noise, including partial translations or raw product identifiers. Therefore, we utilized BLIP-2 [29] to generate new, more descriptive captions.
Dataset Split.
Dataset Cleaning.
In order to mitigate false negatives in our results, we utilized Locality Sensitive Hashing and OpenCLIP ViT-B/16 embeddings to eliminate duplicates. Specifically, we removed duplicates between the test targets and test distractors, as well as between the validation targets and validation distractors. Throughout our experiments, we did not observe any false negatives in the results. However, there remains a small quantity of near-duplicates among the distractor images.
3.2 Composition
In total, we extracted 272,451 products for training, represented in 841,718 images. This represents 581,526 potential simple/complex positive pairs. We additionally extracted 400 products (800 images) to create a validation set, and 2,000 products (4,000 images) for a test set. We added 99,541 simple images in the validation gallery as distractors, and 2,000,014 in the test gallery.
Please refer to Appendix A.4 for metadata details, A.5 for considerations regarding privacy and biases and C for metadata details and a datasheet [13].
3.3 Benchmark
For this dataset, we propose to frame the benchmark as an asymmetric task : the representation of simple images (the gallery) should not be computed conditionally. This choice is motivated by three reasons. First, when using precise free-form conditioning (such as LAION texts, which contain hashed product identifiers and product names) a symmetric encoding would enable a retrieval based solely on this information, completely disregarding the image query. Second, for discrete (categorical) conditioning it allows the presence of items of unknown category in the gallery, which is a situation that may occur in distractors. Third, these images only depict a single object, thus making referring information unnecessary. A similar setting is used by Asai et al. [1].
@@ 4 Conditional Embedding
Figure 4: Overview of our method on LRVS-F. For each element in a batch, we embed the scene conditionally and the isolated item unconditionally. We optimize an InfoNCE loss over the cosine similarity matrix. $\oplus$ denotes concatenation to the patch sequence.
Task Formulation.
@@
We implement $\phi$ by modifying the Vision Transformer (ViT) architecture [9]. The conditioning is an additional input token with an associated learnable positional encoding, concatenated to the sequence of image patches. The content of this token can either be learned directly (e.g. for discrete categorical conditioning), or be generated by another network (e.g. for textual conditioning). At the end of the network, we linearly project the [CLS] token to map the features to a metric space. We experimented with concatenating at different layers in the transformer, and found that concatenating before the first layer is the most sensible choice (see Appendix B.1).
We train the network with the InfoNCE loss [44, 38], following CLIP [40], which is detailed in the next paragraph. However, we hypothesize that even though our method relies on a contrastive loss, it does not explicitly require a specific formulation of it. We choose the InfoNCE loss because of its popularity and scalability. During training, given a batch of $N$ pairs of images and conditioning $((x_{i}^{A},c_{i}^{A});(x_{i}^{B},c_{i}^{B}))_{i=1..N}$, we compute their conditional embeddings $(z_{i}^{A},z_{i}^{B})_{i=1..N}$ with $z=\phi(x,c)\in\mathbb{R}^{d}$. We compute a similarity matrix $S$ where $S_{ij}=s(z_{i}^{A},z_{j}^{B})$, with $s$ the cosine similarity. We then optimize the similarity of the correct pair with a cross-entropy loss, effectively considering the $N-1$ other products in the batch as negatives:
@@
with $\tau$ a learned temperature parameter, and the final loss is $\mathcal{L}=l(S)/2+l(S^{\top})/2$. Please refer to Fig. 4 for an overview of the method. The $\tau$ parameter is used to follow the initial formulation of CLIP [40] and is optimized by gradient during the training. At test time, we use FAISS [22] to create a unique index for the entire gallery and perform fast similarity search on GPUs.